OpenStudy (kyhannah):

Check all that apply. More than one answer. sec 2x= A. 1/cos^2x - sin^2x B. 1/sin^2x - cos^2x C. 1/2sin^2x - 1 D. 1/1-2sin^2x

OpenStudy (anonymous):

abc or d

OpenStudy (kyhannah):

ya it could be one of those

OpenStudy (anonymous):

sorry don't know

OpenStudy (kyhannah):

thats alright :)

OpenStudy (anonymous):

a_d

OpenStudy (anonymous):

a and d I think

OpenStudy (kyhannah):

how did you get a and d M A?

OpenStudy (anonymous):

secx=1/cosx sec2x=1/cos2x cos2x= cos^2-sin^2 cos2x=1-2sin^2

OpenStudy (kyhannah):

ah ok :) Thank you!!!

OpenStudy (anonymous):

yourwelcome

OpenStudy (kyhannah):

:D

OpenStudy (anonymous):

;)